# Learning Equivariant Maps with Variational Quantum Circuits

### v1

@@ Abstract
Geometric quantum machine learning uses the symmetries inherent in data to design tailored machine learning tasks with reduced search space dimension. The field has been well-studied recently in an effort to avoid barren plateau issues while improving the accuracy of quantum machine learning models. This work explores the related problem of learning an equivariant map given two unitary representations of a finite group, which in turn allows the symmetric encoding of the data to be learned rather than simply required. Moreover, this procedure allows the learning of covariant quantum channels, which are an essential tool in quantum information theory. We demonstrate the feasibility of this task and give examples to illustrate the procedure.
December 16, 2024
@@ I Introduction
Quantum machine learning [1, 2, 3, 4] has been a subject of vast interest in recent years for its feasibility in the era of noisy intermediate scale quantum (NISQ) devices [5]. As the name suggests, it aims to combine the success of machine learning with the properties of quantum theory to perform learning in an efficient and accurate manner. While many avenues have been explored, a popular scheme is to make use of a variational quantum circuit in place of a neural network, and for this reason, this circuit is often called a quantum neural network. The parameters are then trained classically through one of the many available optimization routines.
A frequent issue that arises while training variational quantum algorithms is an overly large expressivity, which introduces barren plateaus into the training landscape [6, 7, 8, 9]. However, it has been shown that the natural symmetries in a given problem can be harnessed to alleviate this issue [10, 11, 12] through the use of representation theory. Geometric quantum machine learning, as discussed at length in [13], uses symmetries in this manner. In this framework, the data encoding is represented as an equivariant map between different representations. By building the circuit with only gates which respect the symmetry of the problem, we can be guaranteed that the resulting estimate completely respects this symmetry, thereby reducing the size of the space through which the model must search for a solution. Physicists can liken this process to Lagrangian mechanics, where symmetries in the evolution of a system simplify the corresponding equations of motion. In the same manner, geometric quantum machine learning can reduce the size of the solution space in QML [10, 11, 12]. Symmetry-preserving circuits have also previously proven useful in developing quantum eigensolvers [14, 15, 16, 17].
@@
As representation theory is a natural language for describing the symmetries inherent in a physical system, the methods presented here are of interest to physicists from a variety of subdisciplines. In particular, equivariant maps often take the form of covariant quantum channels [18]. A quantum channel $\mathcal{N}:\mathcal{H}_{A}\to\mathcal{H}_{B}$ mapping between quantum states is said to be covariant with respect to a group $G$ with representations $W_{A}$ and $W_{B}$ if
@@
for all $g\in G$, where $\mathcal{W}_{A}(g)\rho=W_{A}(g)\rho W_{A}(g)^{\dagger}$. Thus, our results facilitate the learning of covariant quantum channels [19, 20]. Physically, these channels describe the symmetric dynamics of a system. For Hamiltonian dynamics, these covariant symmetries can represent forbidden state transitions [21] or superselection rules [22, 23]. In resource theory, covariant channels cannot increase asymmetry measures — a fact which can be thought of as an extension of Noether’s theorem for quantum states [24].
Bijective equivariant maps also form a natural notion of equivalence between representations, and the classification of the representations of a group according to this equivalence relation are of interest to mathematicians working in a variety of subdisciplines. Thus, the ability to learn an equivariant map between equivalent representations is desirable, and the ability to give a measure of how equivariant a map is will be of interest to those studying maps which respect only a substructure between representations.
The work proceeds as follows. In Section II, we review the construction of symmetry-respecting variational quantum circuits and show a simple example of this process in Section III. In Section IV, we show how the methods of the previous sections can be inverted to learn an equivariant map between two given representations, assuming such a map exists. Examples are then provided to illustrate this methodology in Sections IV.1-IV.4. Finally, we give concluding remarks in Section V.
@@ II Symmetry Assisted Quantum Machine Learning
We will begin by reviewing the theory of this problem, restricting ourselves to discrete symmetries for the scope of this work. A unitary representation $W:G\to\mathcal{U}(\mathcal{H})$ of a finite group $G$ is a group homomorphism of $G$ into the group of unitary matrices acting on the Hilbert space $\mathcal{H}$. The unitary condition is necessary in order to implement the represented group element on a quantum device without resorting to block encoding. When $W$ is injective, we call the representation faithful, as its image is isomorphic to $G$ by the first isomorphism theorem [25]. A faithful unitary representation can therefore be thought of as a translation of the description of the group from its abstract group language into the language of quantum gates. Representations that are not faithful respect only a substructure of the original group $G$ in the sense that the image under $W$ is isomorphic to a proper subgroup of $G$.
Given a data set $\mathcal{S}=\{(x_{i},y_{i})\}$ consisting of pairs of data points $x_{i}$ drawn from some data space $\mathcal{R}$ with labels $y_{i}$ in some set $\mathcal{L}$, we assume there is a function $f:\mathcal{R}\to\mathcal{L}$ which correctly maps a data point to its associated label, i.e. $f(x_{i})=y_{i}$ for all $i$. We will make the assumption that $\mathcal{R}$ is endowed with a vector space structure. Our goal is to approximate $f$ using a variational quantum circuit (VQC). Of course, this requires us to embed the data into a Hilbert space [26], and this alone is a nontrivial piece of the VQC, as the choice of embedding can have profound consequences on the trainability of the model (see [27] for robustness results). For now, we will assume that an embedding $\mathcal{E}:\mathcal{R}\to\mathcal{U}(\mathcal{H})$ has been chosen.
@@
$h_{\theta}(x)=\operatorname{Tr}\left[U_{\theta}(\mathcal{E}(x)\ket{0}\!\!\bra{0}\mathcal{E}(x)^{\dagger})U^{\dagger}_{\theta}\mathcal{O}\right],$ (2)
where $\mathcal{O}$ is some chosen observable, and $U_{\theta}$ is the parametrized unitary at the heart of the VQC.
Suppose now that $V:G\to GL(\mathcal{R})$ is a representation of $G$ on the data space. Here and throughout, $GL(\cdot)$ will refer to the general linear group acting on some space. The target function $f$ is called $G$-invariant if
@@
for all $g\in G$ and for all $x\in\mathcal{R}$. This phenomenon appears, for example, in many image classification problems where the label associated to the image does not change if the image is rotated. Since we are trying to estimate $f$, it makes sense to introduce an inductive bias into the model by demanding $G$-invariance in the estimate of $f$. One way to accomplish this is by adding a penalty term to the loss function when the optimization routine is performed which penalizes the output when it lacks $G$-invariance. This property can be efficiently tested for on a quantum computer as in [28]; however, this often extends the training time of the model. An alternative approach is to build the symmetry directly into the VQC, and to do this, we need a little more representation theory.
The representations $V$ and $W$ are called equivariant if there exists a linear transformation $T:\mathcal{R}\to\mathcal{H}$ such that
@@
that is, mapping $\mathcal{R}$ to $\mathcal{H}$ and then applying $W(g)$ is the same as first applying $V(g)$ and then mapping the result to $\mathcal{H}$. When $T$ is bijective, we call it an isomorphism of the representations. If we are given the representation $V:G\to GL(\mathcal{R})$ and a linear map $T:\mathcal{R}\to\mathcal{H}$, then a representation $W:G\to\mathcal{U}(\mathcal{H})$ is said to be induced by $T$ whenever $T$ is an equivariant map between the two representations.
In the context of a VQC, we are given a problem with some symmetry described by a representation $V$. If we are also given an embedding $\mathcal{E}$, then a representation $W$ induced by $\mathcal{E}$ encodes the symmetry of the problem in the Hilbert space picture. In other words, the embedding carries the symmetry structure from the data space to the Hilbert space so that this structure can be respected by the VQC.
An operator $A$ is called $G$-invariant when
@@
for all $g\in G$. That is, $A$ belongs to the commutant of $W(G)$. We now show how to enforce symmetry directly in the VQC by insuring $G$-invariance throughout the circuit.
@@
Let $V$ be a representation of $G$ on the data space $\mathcal{R}$ and let $\mathcal{E}$ be an embedding of the data space into $\mathcal{H}$. Let $W$ be a representation induced by the embedding. Then the estimate $h_{\theta}$ defined by (2) is $G$-invariant if the parameterized unitary $U_{\theta}$ and the observable $\mathcal{O}$ are $G$-invariant. That is,
$h_{\theta}(V(g)x)=h_{\theta}(x)$ (6)
for all $x\in\mathcal{R}$ and $g\in G$ whenever $U_{\theta}$ and $\mathcal{O}$ are $G$-invariant with respect to $W$.
@@
$h_{\theta}(V(g)x)$ (7)
$=\operatorname{Tr}\left[U_{\theta}\mathcal{E}(V(g)x)\ket{0}\!\!\bra{0}\mathcal{E}(V(g)x)^{\dagger}U^{\dagger}_{\theta}\mathcal{O}\right]$ (8)
$=\operatorname{Tr}\left[U_{\theta}W(g)\mathcal{E}(x)\ket{0}\!\!\bra{0}\mathcal{E}(x)^{\dagger}W(g)^{\dagger}U^{\dagger}_{\theta}\mathcal{O}\right].$ (9)
Now applying the $G$-invariance of $U_{\theta}$ produces
$=\operatorname{Tr}\left[W(g)U_{\theta}(\mathcal{E}(x)\ket{0}\!\!\bra{0}\mathcal{E}(x)^{\dagger})U^{\dagger}_{\theta}W(g)^{\dagger}\mathcal{O}\right],$ (10)
@@
$=\operatorname{Tr}\left[U_{\theta}(\mathcal{E}(x)\ket{0}\!\!\bra{0}\mathcal{E}(x)^{\dagger})U^{\dagger}_{\theta}W(g)^{\dagger}\mathcal{O}W(g)\right]$ (11)
$=\operatorname{Tr}\left[U_{\theta}(\mathcal{E}(x)\ket{0}\!\!\bra{0}\mathcal{E}(x)^{\dagger})U^{\dagger}_{\theta}\mathcal{O}\right]$ (12)
$=h_{\theta}(x).$ (13)
Through Proposition 1, we see that that we can force our estimate of $f$ to be $G$-invariant by using only gates and observables which commute with the representation induced by the embedding. The question now is how do we come up with a satisfactory set of such operations? The following lemma is helpful with this problem.
@@
If $[X,Y]=0$, then $[e^{i\theta X},Y]=0$ for all $\theta\in\mathbb{R}$.
@@
$[e^{i\theta X},Y]$ $=\left[\sum_{n=0}^{\infty}\frac{i^{n}\theta^{n}X^{n}}{n!},Y\right]$ (14)
$=\sum_{n=0}^{\infty}\frac{i^{n}\theta^{n}}{n!}[X^{n},Y].$ (15)
@@
By Stone’s theorem [29], for every strongly continuous one-parameter unitary group $U_{\theta}$, there is a self-adjoint operator $H$ such that $U_{\theta}=e^{i\theta H}$, and this is why the lemma is so useful; it tells us that we can find parametrized gates that commute with the representation induced by the embedding by checking that the infinitesimal generator commutes with this representation. For example, if $[Z,W(g)]=0$ for all $g\in G$, then every rotation gate along the $Z$-axis is $G$-invariant.

### v2

@@ Abstract
Geometric quantum machine learning uses the symmetries inherent in data to design tailored machine learning tasks with reduced search space dimension. The field has been well-studied recently in an effort to avoid barren plateau issues while improving the accuracy of quantum machine learning models. This work explores the related problem of learning an equivariant map given two unitary representations of a finite group, which in turn allows the symmetric embedding of the data to be learned rather than simply required. Moreover, this procedure allows the learning of covariant quantum channels, which are an essential tool in quantum information theory. We demonstrate the feasibility of this task and give examples to illustrate the procedure.
February 11, 2025
@@ I Introduction
Quantum machine learning cao2020 ; grant2018 ; schuld2018 ; schuld2020circuit has been a subject of vast interest in recent years for its feasibility in the era of noisy intermediate scale quantum (NISQ) devices preskill2018quantum . It aims to combine the success of machine learning with the properties of quantum theory to perform learning in an efficient and accurate manner. While many avenues have been explored, a popular scheme is to make use of a variational quantum circuit in place of a neural network bishop2006 ; vapnik1999 . The parameters are then trained classically through one of the many available optimization routines.
A frequent issue that arises while training variational quantum algorithms is an overly large expressivity, which introduces barren plateaus into the training landscape arrasmith2022 ; cerezo2021 ; holmes2022 ; mcClean2018 . However, it has been shown that the natural symmetries in a given problem can be harnessed to alleviate this issue larocca2022 ; meyer2023 ; nguyen2022 , and geometric quantum machine learning, as discussed at length in ragone2023 , uses symmetries in this manner. In this framework, the data embedding is represented as an equivariant map between different representations. By building the circuit with only gates which respect the symmetry of the problem, we can be guaranteed that the resulting estimate completely respects this symmetry, thereby reducing the size of the space through which the model must search for a solution. This process mimics the physics of Lagrangian mechanics, where symmetries in the evolution of a system simplify the corresponding equations of motion. In the same manner, geometric quantum machine learning can reduce the size of the solution space in QML larocca2022 ; meyer2023 ; nguyen2022 . Such symmetry-preserving circuits have previously proven useful in developing quantum eigensolvers SSY20 ; Gard2020 ; BGAMBE21 ; LXYB22 .
@@
As representation theory is a natural language for describing the symmetries inherent in a physical system, the methods presented here are of interest to physicists from a variety of subdisciplines. In particular, equivariant maps often take the form of covariant quantum channels holevo2002 . A quantum channel $\mathcal{N}:\mathcal{H}_{A}\to\mathcal{H}_{B}$ mapping between quantum states is said to be covariant with respect to a group $G$ with representations $W_{A}$ and $W_{B}$ if
@@
for all $g\in G$, where $\mathcal{W}_{A}(g)\rho=W_{A}(g)\rho W_{A}(g)^{\dagger}$. Thus, our results facilitate the learning of covariant quantum channels CDP09 ; cemin2024machinelearningquantumchannels . Physically, these channels describe the symmetric dynamics of a system. For Hamiltonian dynamics, these covariant symmetries can represent forbidden state transitions vojta2003quantum or superselection rules Wick1952 ; Aharonov1967superselection . In resource theory, covariant channels cannot increase asymmetry measures — a fact which can be thought of as an extension of Noether’s theorem for quantum states marvian2014Noether .
Bijective linear equivariant maps, sometimes referred to as intertwining maps, also form a natural notion of equivalence between representations hall2003 ; steinberg2011 , and the classification of the representations of a group according to this equivalence relation are of interest to mathematicians working in a variety of subdisciplines such as geometry, statistics, and representation theory itself. Thus, the ability to learn an intertwining map between equivalent representations is desirable, and the ability to give a measure of how equivariant a map is will be of interest to those studying maps which respect only a substructure between representations.
Deep learning theorists also rely on equivariant maps to utilize data symmetries for improved performance even in classical machine learning mondal2023advances ; ouderaa2023learning ; minartz2023neuralsimulators ; wang2022approximately . Graph neural networks, for instance, must respect permutation symmetry, making learning equivariance an important task in that framework huang2023approx . The task of learning equivariant maps for classical deep neural networks has been recently investigated in ouderaa2023learning ; benton2020learning , and naturally arises when considering symmetry-constrained dynamics wang2022approximately .
The rest of this work is structured as follows: In Section II, we review the construction of symmetry-respecting variational quantum circuits, and we complement this theoretical discussion with a concrete example in Section II.1, where we solve the toy problem of classifying vertical and horizontal lines in $2\times 2$ images with a single parameter in a variational quantum machine learning algorithm that respects the geometric structure of this problem. In Section III, we show how the methods of the previous section can be inverted to learn an equivariant map between two given representations, assuming such a map exists. Several examples are then provided to illustrate this methodology. In Section IV, we show that this method can be extended to learning intertwining maps, despite the data embedding being nonlinear. Finally, we give concluding remarks in Section V.
@@ II Symmetry Assisted Quantum Machine Learning
We begin by reviewing the mathematical foundations of this problem, restricting our attention to discrete symmetries for the scope of this work. A unitary representation $W:G\to\mathcal{U}(\mathcal{H})$ of a finite group $G$ is a group homomorphism of $G$ into the group of unitary matrices acting on the Hilbert space $\mathcal{H}$. The unitary condition is assumed so that quantum states are mapped to quantum states. When $W$ is injective, we call the representation faithful, as its image is isomorphic to $G$ by the first isomorphism theorem (dummit2004, , Chapter 3). Such a representation can therefore be thought of as a translation of the description of the group from its abstract group language into the language of quantum gates. Representations that are not faithful respect only a substructure of the original group $G$ in the sense that the image under $W$ is isomorphic to a proper subgroup of $G$.
We aim to apply representation theory to the study of classification problems. Given a data set $\mathcal{S}=\{(x,y)\}$ consisting of pairs of data points $x$ drawn from some data space $\mathcal{R}$ with labels $y$ in some set $\mathcal{L}$, we assume there is a function $f:\mathcal{R}\to\mathcal{L}$ which correctly maps a data point to its associated label, i.e. $f(x)=y$. We will make the assumption that $\mathcal{R}$ is endowed with a vector space structure. Typically, we approximate $f$ using a variational quantum circuit (VQC). Of course, this requires us to embed the data into a Hilbert space lloyd2020quantum , and this alone is a nontrivial piece of the VQC, as the choice of embedding can have profound consequences on the trainability of the model (see laRose2020 for robustness results). For now, we will assume that an embedding $\mathcal{E}:\mathcal{R}\to\mathcal{U}(\mathcal{H})$ has been chosen, which by convention we always take to act on the initial all-zeroes state, |$\ket{\vec{0}}$, which could in principle consist of multiple qubits.
@@
$f_{\phi}(x)=\operatorname{Tr}\left[U_{\phi}(\mathcal{E}(x)\ket{\vec{0}}\!\!\bra{\vec{0}}\mathcal{E}(x)^{\dagger})U^{\dagger}_{\phi}\mathcal{O}\right],$ (2)
where $\mathcal{O}$ is some chosen observable, and $U_{\phi}$ is the parametrized unitary at the heart of the VQC. In QML, $f_{\phi}$ is then used to model the behavior of $f$ as closely as possible by ensuring that $\|f_{\phi}(x)-y\|$ is minimal for every data-label pair $(x,y)$.
Suppose now that the data respects some symmetry described by a group $G$ with representation $V:G\to GL(\mathcal{R})$, where $GL(\mathcal{R})$ is the general linear group on $\mathcal{R}$. The target function $f$ is called $G$-invariant if
@@
for all $g\in G$ and for all $x\in\mathcal{R}$. This phenomenon appears, for example, in many image classification problems where the label associated to the image does not change if the image is rotated.
Since we are trying to estimate $f$ with $f_{\phi}$, it makes sense to introduce an inductive bias into the VQC model by demanding $G$-invariance. One way to accomplish this is by adding a penalty term to the loss function when the optimization routine is performed which penalizes the output when it lacks $G$-invariance. This property can be efficiently tested for on a quantum computer as in laBorde2023 ; however, this often extends the training time of the model. An alternative approach is to build the symmetry directly into the VQC, and to do this, we need a notion of equivariance.
The representations $V$ and $W$ are called equivariant if there exists a map $T:\mathcal{R}\to\mathcal{H}$ (called an equivariant map) such that
@@
which bears a striking resemblance to the covariance condition given in (1). When $T$ is bijective and linear, we call it an isomorphism of the representations, and the representations themselves are called equivalent. If we are given the representation $V:G\to GL(\mathcal{R})$ and a map $T:\mathcal{R}\to\mathcal{H}$, then a representation $W:G\to\mathcal{U}(\mathcal{H})$ is said to be induced by $T$ whenever $T$ is equivariant.
Additionally, an operator $A$ is called $G$-invariant when
@@
for all $g\in G$. That is, $A$ belongs to the commutant of $W(G)$. This fact is important in the coming derivations, as we enforce symmetry directly in the VQC by ensuring $G$-invariance throughout the circuit.
In the context of a VQC, we are given a problem with some symmetry described by a representation $V$. If we are also given an embedding $\mathcal{E}$, then a representation $W$ induced by $\mathcal{E}$ encodes the symmetry of the problem in the Hilbert space picture. In other words, the embedding carries the symmetry structure from the data space to the Hilbert space so that this structure can be respected by the VQC. Strictly speaking, the embedding is a map $\mathcal{E}:\mathcal{R}\to\mathcal{U}(\mathcal{H})$ which by convention acts on the all-zeroes state |$\ket{\vec{0}}$, thus the map $T$ in (4) is really given by $T(x)=\mathcal{E}(x)\ket{\vec{0}}$. When we say that $W$ is induced by the embedding, what we really mean is that $W$ is induced by the map $T:\mathcal{R}\to\mathcal{H}$, which is defined by the embedding. If instead we are given two representations and we want to guarantee that the embedding is an equivariant map between them, then this is a non-trivial task which we take up in Section III.
@@
Let $V$ be a representation of $G$ on the data space $\mathcal{R}$ and let $\mathcal{E}$ be an embedding of the data space into $\mathcal{H}$. Let $W$ be a representation induced by the embedding. Then the estimate $f_{\phi}$ defined by (2) is $G$-invariant if the parameterized unitary $U_{\phi}$ and the observable $\mathcal{O}$ are $G$-invariant. That is,
$f_{\phi}(V(g)x)=f_{\phi}(x)$ (6)
for all $x\in\mathcal{R}$ and $g\in G$ whenever $U_{\phi}$ and $\mathcal{O}$ are $G$-invariant with respect to $W$.
@@
$f_{\phi}(V(g)x)$ $=\operatorname{Tr}\left[U_{\phi}\mathcal{E}(V(g)x)\ket{0}\!\!\bra{0}\mathcal{E}(V(g)x)^{\dagger}U^{\dagger}_{\phi}\mathcal{O}\right]$ (7)
$=\operatorname{Tr}\left[U_{\phi}W(g)\mathcal{E}(x)\ket{0}\!\!\bra{0}\mathcal{E}(x)^{\dagger}W(g)^{\dagger}U^{\dagger}_{\phi}\mathcal{O}\right].$ (8)
Now applying the $G$-invariance of $U_{\phi}$ produces
$=\operatorname{Tr}\left[W(g)U_{\phi}\mathcal{E}(x)\ket{0}\!\!\bra{0}(\mathcal{E}(x))^{\dagger}U^{\dagger}_{\phi}(W(g))^{\dagger}\mathcal{O}\right],$ (9)
@@
$=\operatorname{Tr}\left[U_{\phi}\mathcal{E}(x)\ket{0}\!\!\bra{0}(\mathcal{E}(x))^{\dagger}U^{\dagger}_{\phi}(W(g))^{\dagger}\mathcal{O}W(g)\right]$ (10)
$=\operatorname{Tr}\left[U_{\phi}\mathcal{E}(x)\ket{0}\!\!\bra{0}(\mathcal{E}(x))^{\dagger}U^{\dagger}_{\phi}\mathcal{O}\right]$ (11)
$=f_{\phi}(x).$ (12)
This completes the proof. ∎
Through Proposition 1, we see that we may enforce $G$-invariance in $f_{\phi}$ by using only gates and observables which commute with the representation induced by the embedding, but how do we come up with a satisfactory set of such operations? The following lemma is helpful.
@@
If $[X,Y]=0$, then $[e^{i\phi X},Y]=0$ for all $\phi\in\mathbb{R}$.
@@
$[e^{i\phi X},Y]$ $=\left[\sum_{n=0}^{\infty}\frac{i^{n}\phi^{n}X^{n}}{n!},Y\right]=\sum_{n=0}^{\infty}\frac{i^{n}\phi^{n}}{n!}[X^{n},Y].$ (13)
@@
By Stone’s theorem hall2013quantum , for every strongly continuous one-parameter unitary group $U_{\phi}$, there is a self-adjoint operator $H$ such that $U_{\phi}=e^{i\phi H}$. This is why the lemma is so useful; it tells us that we can find parametrized gates that commute with the representation induced by the embedding by checking that the infinitesimal generator commutes with this representation. For example, if $[Z,W(g)]=0$ for all $g\in G$, then every rotation gate along the $Z$-axis is $G$-invariant.
@@
and this completes the proof. ∎
This lemma gives a technique for constructing VQCs which respect the symmetry of a problem. First find a representation $W$ induced by the embedding $\mathcal{E}$. Then construct a $G$-invariant gate set by twirling self-adjoint operators and using them as the infinitesimal generators of the gate set. Finally, choose an observable which is $G$-invariant. By Proposition 1, the estimate produced by the model will now be $G$-invariant. Given any three of these building blocks —the embedding, the initial group representation, the embedded representation, and the observable measured— the remaining piece could in principle be learned or determined to suit the task at hand, and we will return to this point in Section III.
II.1 Example: Classifying Lines in Images
Let us demonstrate how to construct VQCs which respect the symmetry of a problem through a simple example. Consider the problem of classifying vertical and horizontal lines in $2\times 2$ grayscale images (see Figure 1) which are easily generated in Python.